# Training Large Language Models to Reason In Parallel with Global Forking Tokens

### v2

@@ A.1 Algorithm: SSFT Implementation
1: $\pi_{{\bm{\theta}}}$: base model $\mathrm{N}$: Number of global forking tokens $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ $\mathcal{D}$: Dataset with (at most) $\mathrm{M}$ reasoning traces per question $\mathrm{B}$: Global batch size $\mathrm{T}_{\mathrm{L}}$}: The first $\mathrm{T}_{\mathrm{L}}$} number of tokens to match in matching cost (Equation  7).
@@
13:   Compute optimal matching $\hat{\bm{\sigma}}_{k}$ between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ and $\{{\mathbf{r}}_{k}^{(j)}\}_{j=1}^{\mathrm{M}}$. Hungarian algorithm (hungarian) can be applied to $\mathrm{C}$ to efficiently compute Equation 9 (Equation 3).
@@
1: $\pi_{{\bm{\theta}}}$: base model $\mathrm{N}$: Number of global forking tokens $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ $\mathcal{D}$: Dataset with (at most) $\mathrm{M}$ reasoning traces per question, variable $m$ number of traces per question $\mathrm{B}$: Global batch size $\mathrm{T}_{\mathrm{L}}$}: The first $\mathrm{T}_{\mathrm{L}}$} number of tokens to match in matching cost (Equation  7). $b$: Original per-device global batch size ($b\geq\mathrm{M}$), This is “micro batch size*original grad accumulation steps”
@@
Algorithm 1 presents the core SSFT implementation with optimal bipartite matching. In practice, the nested-loop computation used to populate $\mathrm{C}$ is fully vectorized and can be executed in a single forward pass. This does not blow up VRAM because (i) we do not store activations for these cost evaluations (no backprop through matching costs), and (ii) We only need to use the first ${\mathrm{T}_{\mathrm{L}}}<{\mathrm{T}_{{\mathbf{r}}}}$ NTP losses to compute the matching cost, as the NTP loss over the first few thousand tokens can already differentiate many reasoning traces in terms of their modes. Nevertheless, our code also supports matching over the full $\mathrm{T}_{{\mathbf{r}}}$} tokens by chunking the computation into a few batches, so this step does not become a VRAM bottleneck. Fine-tuning on 1k questions with 4 traces each, SSFT (optimal matching) took 6.5 h for 6 epochs, compared to 6.1 h for standard SFT, adding only $6.6\%$ overhead.
The primary VRAM bottleneck in SSFT remains the backpropagation Step 10, regardless of whether we use optimal or random matching, because the effective batch size scales with $\mathrm{M}$. To mitigate this, we split the backward pass into several gradient-accumulation steps. Although our experiments use the same number of reasoning traces per question, we also support variable number of targets using our queue-based batching in Algorithm 2. The complication arises when using distributed training with a variable-sized batch, as different processes require the same per-device batch size to perform collective operations. This is mitigated by padding with “PAD sequences” to align batch sizes. Our implementation minimizes the number of “PAD sequences” by storing a variable number of targets in a queue and dequeuing multiple items to form a per-device global batch, so smaller batches can be stitched together instead of always being padded.
A.2 Learned Matchings between Global Forking Tokens and Traces
A.2.1 How to Choose the Global Forking Token for Pass@1
A.3 Training details
A.3.1 Training Datasets
32B experiments with questions from s1(Main): We explain the process of generating our training dataset for experiments in Table 1, Figure 4, Figure 5, Figure 6, Figure 7. First, we use the 1000 questions from s1 (muennighoff2025s1_simpletesttime) and populate a pool of reasoning traces by distilling from GPT-OSS-120B-high reasoning, GPT-OSS-120B-medium reasoning, DeepSeek R1, Gemini Flash2.0 Thinking, and Claude4/4.1. We use temperature 1.0, maximum length of 32768, and sets high reasoning effort unless specified. We generate two traces per teacher model. We use Claude3.5 Sonnet to extract the answer from the distilled solutions and compare with the ground-truth answer. The correctness of these distilled traces are shown in 3.
@@ A.3.1 Training Datasets
A.3.2 Training Hyperparameters
For consistency, we use Qwen2.5-32B-Instruct (yang2025qwen3) as the base model for all of our 32B experiments. We use standard fine-tuning hyperparameters: we train for 6 epochs with a global batch size of 32, which is derived from 4 gradient accumulation steps and distributed training with 8 GPUS ($4\times 8=32$). This results in 756 gradient steps. The maximum sequence length is set to 32,768. We train with bfloat16 with a learning rate of 1e51e-5 warmed up linearly for 5% and then decayed to 0 using a cosine schedule. We choose AdamW optimizer (loshchilov2017decoupled) with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and weight decay 1e41e-4. We only backpropagates the completion loss, which is the loss on reasoning traces and the answers. Fine-tuning SSFT-32B plus loggings took 6.5 hours on 8 NVIDIA B200 GPUs using PyTorch FSDP, Liger Kernel (hsu2024ligerkernel) for fused cross entropy loss, and FlashAttention-2 (dao2023flashattention) for fused attention computation. Fine-tuning SSFT-32B (random $\bm{\sigma}$) took 6.3 hours, and Fine-tuning SFT-mixed-distill-32B took 6.1 hours. Even our baseline SFT-OSS-distill-32B with only one trace per question, and our attempt to reproduce s1.1 took 1.66 hours, which is longer than the time reported by muennighoff2025s1_simpletesttime. This is due to using 8 GPUs instead of 16 GPUs, hardware and package differences. When training with s1k-4mixed-reasoning, we added one extra epoch from 5 epochs to 6 epochs, since we have 4x reasoning traces, but we did not linearly increase the number of epochs, as these traces can be similar, and the number of distinct questions is still 1,000. Overall, we made sure all of our models are fine-tuned with consistent hyperapameters.
A.3.3 Visualization of SSFT Training Dynamics
@@
A.3.4 Ablation Study Training details (Removing High Quality Small Dataset)
For this ablation study, we choose Open-R1-Math220k default split, which has 93,000 math questions and $2\sim 4$ traces. Since Qwen2.5-Math-7B is a widely fine-tuned model using this dataset, we also choose it as our base model. We train for 3 epochs using 8 A100 GPUs, which took around 4 days. Our hyperparameters are mostly consistent with the recommended hyperparameters by openr1. We fine-tune both SSFT-OpenR1-93k-7B and SFT-OpenR1-93k-7B with a maximum length of 32768, learning rate of 4.0e054.0e-05 warmed up linearly for $3\%$ and decayed to 0 following cosine schedule, 8 gradient accumulation steps. For our SSFT method, we reserve $\mathrm{N}=4$ global forking tokens, and use the first 1000 tokens for matching. Again, only the completion loss is used for optimizing the model parameters.
A.4 Examples of SSFT-32B Reasoning Traces initiated by Different Global Forking Tokens
@@
A.5 Alternative Implementation of training with multi-target sequences
Figure 11 shows an alternative implementation for training with multiple target parallel sequences. Instead of flattening multi-target data along the batch dimension, we can also concatenate parallel reasoning traces along the sequence dimension, modify the causal attention matrix, and position ids for positional embeddings. While easy to implement, this version extends the sequence length and cannot perform gradient accumulation along the batch dimension to maintain the same VRAM as training with a single reasoning trace. Since most fine-tuning already sets the micro-batch size to 1 due to the increased number of reasoning tokens already present in a single reasoning path, this implementation also cannot further shrink the micro-batch size to accommodate the extended context length. In addition, the code that uses the initial flash-attention implementation, which requires causal attention matrix, is not compatible with this setup. This implementation is also less flexible in maintaining a consistent number of sequences per global batch.

### v3

@@ A.8 Algorithm: SSFT Implementation
1: $\pi_{{\bm{\theta}}}$: base model $\mathrm{N}$: Number of global forking tokens $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ $\mathcal{D}$: Dataset with (at most) $\mathrm{M}$ reasoning traces per question $\mathrm{B}$: Global batch size $\mathrm{T}_{\mathrm{L}}$}: The first $\mathrm{T}_{\mathrm{L}}$} number of tokens to match in matching cost (Equation  5).
@@
13:   Compute optimal matching $\hat{\bm{\sigma}}_{k}$ between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ and $\{{\mathbf{r}}_{k}^{(j)}\}_{j=1}^{\mathrm{M}}$. Hungarian algorithm (Kuhn, 1955) can be applied to $\mathrm{C}$ to efficiently compute Equation 7 (Equation 1).
@@
1: $\pi_{{\bm{\theta}}}$: base model $\mathrm{N}$: Number of global forking tokens $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ $\mathcal{D}$: Dataset with (at most) $\mathrm{M}$ reasoning traces per question, variable $m$ number of traces per question $\mathrm{B}$: Global batch size $\mathrm{T}_{\mathrm{L}}$}: The first $\mathrm{T}_{\mathrm{L}}$} number of tokens to match in matching cost (Equation  5). $b$: Original per-device global batch size ($b\geq\mathrm{M}$), This is “micro batch size*original grad accumulation steps”
@@
A.9 Training details
A.9.1 Training Datasets
32B experiments with questions from s1(Main): We explain the process of generating our training dataset for experiments in Table 1, Figure 3, Figure 4, Figure 5, Figure 6. First, we use the 1000 questions from s1 (Muennighoff et al., 2025) and populate a pool of reasoning traces by distilling from GPT-OSS-120B-high reasoning, GPT-OSS-120B-medium reasoning, DeepSeek R1, Gemini Flash2.0 Thinking, and Claude4/4.1. We use temperature 1.0, maximum length of 32768, and sets high reasoning effort unless specified. We generate two traces per teacher model. We use Claude3.5 Sonnet to extract the answer from the distilled solutions and compare with the ground-truth answer. The correctness of these distilled traces are shown in 8.
@@ A.9.1 Training Datasets
A.9.2 SSFT Training Details
For consistency, we use Qwen2.5-32B-Instruct (Yang et al., 2025a) as the base model for all of our 32B experiments. We use standard fine-tuning hyperparameters: we train for 6 epochs with a global batch size of 32, which is derived from 4 gradient accumulation steps and distributed training with 8 GPUS ($4\times 8=32$). This results in 756 gradient steps. The maximum sequence length is set to 32,768. We train with bfloat16 with a learning rate of 1e51e-5 warmed up linearly for 5% and then decayed to 0 using a cosine schedule. We choose AdamW optimizer (Loshchilov and Hutter, 2017) with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and weight decay 1e41e-4. We only backpropagates the completion loss, which is the loss on reasoning traces and the answers. Fine-tuning SSFT-32B plus loggings took 6.5 hours on 8 NVIDIA B200 GPUs using PyTorch FSDP, Liger Kernel (Hsu et al., 2024) for fused cross entropy loss, and FlashAttention-2 (Dao, 2023) for fused attention computation. Fine-tuning SSFT-32B (random $\bm{\sigma}$) took 6.3 hours, and Fine-tuning SFT-mixed-distill-32B took 6.1 hours. Even our baseline SFT-OSS-distill-32B with only one trace per question, and our attempt to reproduce s1.1 took 1.66 hours, which is longer than the time reported by Muennighoff et al. (2025). This is due to using 8 GPUs instead of 16 GPUs, hardware and package differences. When training with s1k-4mixed-reasoning, we added one extra epoch from 5 epochs to 6 epochs, since we have 4x reasoning traces, but we did not linearly increase the number of epochs, as these traces can be similar, and the number of distinct questions is still 1,000. Overall, we made sure all of our models are fine-tuned with consistent hyperapameters.
A.9.3 Visualization of SSFT Training Dynamics
@@
A.9.4 Ablation Study Training details (Removing High Quality Small Dataset)
For this ablation study, we choose Open-R1-Math220k default split, which has 93,000 math questions and $2\sim 4$ traces. Since Qwen2.5-Math-7B is a widely fine-tuned model using this dataset, we also choose it as our base model. We train for 3 epochs using 8 A100 GPUs, which took around 4 days. Our hyperparameters are mostly consistent with the recommended hyperparameters by Face (2025). We fine-tune both SSFT-OpenR1-93k-7B and SFT-OpenR1-93k-7B with a maximum length of 32768, learning rate of 4.0e054.0e-05 warmed up linearly for $3\%$ and decayed to 0 following cosine schedule, 8 gradient accumulation steps. For our SSFT method, we reserve $\mathrm{N}=4$ global forking tokens, and use the first 1000 tokens for matching. Again, only the completion loss is used for optimizing the model parameters.
A.9.5 Other Training Setup and Hyperparameters
Table 9 shows the hyperparameters used for all of our key experiments.
Table 9: Trainning details of SSFT and GFPO. SFT-mixed-distill-32B-tags-GRPO optimizes over full sequences, whereas GFPO only optimizes <think i> , tokenized in 4 tokens. 8xB200 is used for full GRPO.
SSFT GFPO (GRPO on Forking Tokens)
Parameter Value Parameter Value
Global Batch Size 32 Global Batch Size 128
GPUs 8xB200 GPUs 8xA100
Max Sequence Length 32768 Max Generation Length 31000
$\mathrm{N}$ Global Forking Tokens 6 Max Prompt Length 1024
$\mathrm{M}$ distillation targets 4 Number of rollouts 5
$\mathrm{L}$ Matching Length 1000 PPO minibatch size 128
Learning Rate 1e-5 Learning Rate 1e-6
Adam $\beta_{1}$ 0.9 Adam $\beta_{1}$ 0.9
Adam $\beta_{2}$ 0.95 Adam $\beta_{2}$ 0.95
Warmup ratio $5\%$ Rollout Temperature 0.7
Epochs 66 Epochs 1
KL coeffcient 0.001
Training data for math models
s1k-4mixed-reasoning s1 questions with 4 mixed traces each. 1280 questions from DAPO-Math-17k
Training data for coding models
code1k-4mixed-reasoning - 1k questions from Open-Thoughts - 4 R1 traces each 1280 questions from Intellect-2-RL
Base models being used
Qwen2.5-32B-Instruct Qwen3-4B-Base Llama3.1-8B-Instruct SSFT-32B SFT-mixed-distill-32B-tags SSFT-32B (random $\bm{\sigma}$) SSFT-32B-code (before RL) SFT-mixed-distill-32B-code (before RL) SSFT-32B (random $\bm{\sigma}$)-code (before RL)
Qwen2.5-Math-7B-Instruct. Separate training details and hyperparameters reported in Appendix A.9.4
A.10 Examples of SSFT-32B Reasoning Traces initiated by Different Global Forking Tokens
@@
A.11 Alternative Implementation of training with multi-target sequences
Figure 11 shows an alternative implementation for training with multiple target parallel sequences. Instead of flattening multi-target data along the batch dimension, we can also concatenate parallel reasoning traces along the sequence dimension, modify the causal attention matrix, and position ids for positional embeddings. While easy to implement, this version extends the sequence length and cannot perform gradient accumulation along the batch dimension to maintain the same VRAM as training with a single reasoning trace. Since most fine-tuning already sets the micro-batch size to 1 due to the increased number of reasoning tokens already present in a single reasoning path, this implementation also cannot further shrink the micro-batch size to accommodate the extended context length. In addition, the code that uses the initial flash-attention implementation, which requires causal attention matrix, is not compatible with this setup. In terms of choosing the number of sequences in a global batch size, this implementation is also less flexible.